OpenStudy (anonymous):

Is 10000058 divisible by 10

OpenStudy (anonymous):

10000058/10=1000005.8

OpenStudy (bookworm14):

What do you think? :)

OpenStudy (anonymous):

so no not evenly at least

OpenStudy (anonymous):

exlplain here are the answer choices No Yes because the last digit is an even number Yes because of the frist two digits of the number are 10

OpenStudy (bookworm14):

The only time a number will divide evenly by 10 is when it ends on a zero examples: 150/10= 15 ..... 370/10=37 ...etc If it does not end in zero, then it will NOT evenly divide examples: 157/10=15.7 ....... 463/10=46.3 ......etc Hope this helps! ~Bree

OpenStudy (anonymous):

is it A B OR C

OpenStudy (bookworm14):

Using the info I gave you, what do you think the answer is (give a guess, then i will tell you the correct answer)

OpenStudy (anonymous):

OKAY I SUBMITTED IT IS b

OpenStudy (bookworm14):

It is A, because no it is not divisible by 10

OpenStudy (bookworm14):

THE ANSWER IS A

OpenStudy (anonymous):

oh ok